OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-3,-6) and (-8,-4)

OpenStudy (dumbcow):

first find slope m = (-6-(-4))/(-3-(-8)) = -2/5 pick a point to use for x,y value to find b -6 = (-2/5)(-3) +b b = -6 - 6/5 = -36/5 y = -2/5 x - 36/5